### v2

M. Ablikim1, M. N. Achasov4,c, P. Adlarson75, O. Afedulidis3, X. C. Ai80, R. Aliberti35, A. Amoroso74A,74C, Q. An71,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao63, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi74A,74C, E. Bianco74A,74C, A. Bortone74A,74C, I. Boyko36, R. A. Briere5, A. Brueggemann68, H. Cai76, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,63, N. Cao1,63, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,63, H. Y. Chen20, M. L. Chen1,58,63, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,63, X. R. Chen31,63, X. T. Chen1,63, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,63, S. K. Choi10A, G. Cibinetto29A, F. Cossio74C, J. J. Cui50, H. L. Dai1,58, J. P. Dai78, A. Dbeyssi18, R. E. de Boer3, D. Dedovich36, C. Q. Deng72, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis74A,74C, F. De Mori74A,74C, B. Ding66,1, X. X. Ding46,h, Y. Ding40, Y. Ding34, J. Dong1,58, L. Y. Dong1,63, M. Y. Dong1,58,63, X. Dong76, M. C. Du1, S. X. Du80, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang1,58, J. Fang59, S. S. Fang1,63, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava74B,74C, F. Feldbauer3, G. Felici28A, C. Q. Feng71,58, J. H. Feng59, Y. T. Feng71,58, M. Fritsch3, C. D. Fu1, J. L. Fu63, Y. W. Fu1,63, H. Gao63, X. B. Gao41, Y. N. Gao46,h, Yang Gao71,58, S. Garbolino74C, I. Garzia29A,29B, L. Ge80, P. T. Ge76, Z. W. Ge42, C. Geng59, E. M. Gersabeck67, A. Gilman69, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco74A,74C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,63, A. Q. Guo31,63, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han63, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris65, K. K. He55, K. L. He1,63, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,63, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,63, X. T. Hou1,63, Y. R. Hou63, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,63, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,63, Y. Hu1, G. S. Huang71,58, K. X. Huang59, L. Q. Huang31,63, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain73, F. Hölzken3, N. Hüsken35, N. in der Wiesche68, J. Jackson27, S. Janchiv32, J. H. Jeong10A, Q. Ji1, Q. P. Ji19, W. Ji1,63, X. B. Ji1,63, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia71,58, D. Jiang1,63, H. B. Jiang76, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,63, Y. Jiang63, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin66, M. Q. Jing1,63, X. M. Jing63, T. Johansson75, S. Kabana33, N. Kalantar-Nayestanaki64, X. L. Kang9, X. S. Kang40, M. Kavatsyuk64, B. C. Ke80, V. Khachatryan27, A. Khoukaz68, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,63, N. Kumar26, A. Kupsc44,75, W. Kühn37, J. J. Lane67, P. Larin18, L. Lavezzi74A,74C, T. T. Lei71,58, Z. H. Lei71,58, M. Lellmann35, T. Lenz35, C. Li43, C. Li47, C. H. Li39, Cheng Li71,58, D. M. Li80, F. Li1,58, G. Li1, H. B. Li1,63, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, L. J. Li1,63, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,63, Q. X. Li50, R. Li17,31, S. X. Li12, T. Li50, W. D. Li1,63, W. G. Li1,a, X. Li1,63, X. H. Li71,58, X. L. Li50, X. Y. Li1,63, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li78, C. Liang42, H. Liang71,58, H. Liang1,63, Y. F. Liang54, Y. T. Liang31,63, G. R. Liao14, L. Z. Liao50, Y. P. Liao1,63, J. Libby26, A. Limphirat60, C. C. Lin55, D. X. Lin31,63, T. Lin1, B. J. Liu1, B. X. Liu76, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,63, Huihui Liu21, J. B. Liu71,58, J. Y. Liu1,63, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu71,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu63, S. B. Liu71,58, T. Liu12,g, W. K. Liu43, W. M. Liu71,58, X. Liu38,k,l, X. Liu39, Y. Liu38,k,l, Y. Liu80, Y. B. Liu43, Z. A. Liu1,58,63, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,63, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,63, C. L. Luo41, J. R. Luo59, M. X. Luo79, T. Luo12,g, X. L. Luo1,58, X. R. Lyu63, Y. F. Lyu43, F. C. Ma40, H. Ma78, H. L. Ma1, J. L. Ma1,63, L. L. Ma50, M. M. Ma1,63, Q. M. Ma1, R. Q. Ma1,63, T. Ma71,58, X. T. Ma1,63, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora74A,74C, S. Malde69, Y. J. Mao46,h, Z. P. Mao1, S. Marcello74A,74C, Z. X. Meng66, J. G. Messchendorp13,64, G. Mezzadri29A, H. Miao1,63, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,63, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen63, Q. Ouyang1,58,63, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A. Pathak34, P. Patteri28A, Y. P. Pei71,58, M. Pelizaeus3, H. P. Peng71,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,63, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi71,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian63, C. F. Qiao63, X. K. Qiao80, J. J. Qin72, L. Q. Qin14, L. Y. Qin71,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu72, C. F. Redmer35, K. J. Ren39, A. Rivetti74C, M. Rolo74C, G. Rong1,63, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning75, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan71,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,63, M. Shao71,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen63, X. Y. Shen1,63, B. A. Shi63, H. Shi71,58, H. C. Shi71,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi72, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y. J. Song12,g, Y. X. Song46,h,n, S. Sosio74A,74C, S. Spataro74A,74C, F. Stieler35, Y. J. Su63, G. B. Sun76, G. X. Sun1, H. Sun63, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun76, S. S. Sun1,63, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun71,58, Y. Z. Sun1, Z. Q. Sun1,63, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang71,58, Y. A. Tang76, L. Y. Tao72, Q. T. Tao25,i, M. Tat69, J. X. Teng71,58, V. Thoren75, W. H. Tian59, Y. Tian31,63, Z. F. Tian76, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang63, Bo Wang71,58, D. Y. Wang46,h, F. Wang72, H. J. Wang38,k,l, J. J. Wang76, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang63, S. Wang12,g, S. Wang38,k,l, T. Wang12,g, T. J. Wang43, W. Wang59, W. Wang72, W. P. Wang35,71,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,63, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L. Wang72, Z. Y. Wang1,63, Ziyi Wang63, D. H. Wei14, F. Weidner68, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson69, M. Wolke75, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,63, X. Wu12,g, X. H. Wu34, Y. Wu71,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia71,58, X. M. Xian39, B. H. Xiang1,63, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y. L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie71,58, T. Y. Xing1,63, C. F. Xu1,63, C. J. Xu59, G. F. Xu1, H. Y. Xu66,2,p, M. Xu71,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu66, X. P. Xu55, Y. C. Xu77, Z. P. Xu42, Z. S. Xu63, F. Yan12,g, L. Yan12,g, W. B. Yan71,58, W. C. Yan80, X. Q. Yan1, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,63, Y. F. Yang43, Y. X. Yang1,63, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Z. Y. You59, B. X. Yu1,58,63, C. X. Yu43, G. Yu1,63, J. S. Yu25,i, T. Yu72, X. D. Yu46,h, Y. C. Yu80, C. Z. Yuan1,63, J. Yuan34, J. Yuan45, L. Yuan2, S. C. Yuan1,63, Y. Yuan1,63, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar73, F. R. Zeng50, S. H. Zeng72, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng1,63, Y. J. Zeng59, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,63, B. L. Zhang1,63, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang80, H. Zhang71,58, H. C. Zhang1,58,63, H. H. Zhang34, H. H. Zhang59, H. Q. Zhang1,58,63, H. R. Zhang71,58, H. Y. Zhang1,58, J. Zhang80, J. Zhang59, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,63, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,63, Jianyu Zhang63, L. M. Zhang61, Lei Zhang42, P. Zhang1,63, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,63, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y. Zhang50, Y. Zhang72, Y. Zhang1, Y. T. Zhang80, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang71,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang76, Z. Y. Zhang43, Z. Z. Zhang45, G. Zhao1, J. Y. Zhao1,63, J. Z. Zhao1,58, L. Zhao1, Lei Zhao71,58, M. G. Zhao43, N. Zhao78, R. P. Zhao63, S. J. Zhao80, Y. B. Zhao1,58, Y. X. Zhao31,63, Z. G. Zhao71,58, A. Zhemchugov36,b, B. Zheng72, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,63, Y. H. Zheng63, B. Zhong41, X. Zhong59, H. Zhou50, J. Y. Zhou34, L. P. Zhou1,63, S. Zhou6, X. Zhou76, X. K. Zhou6, X. R. Zhou71,58, X. Y. Zhou39, Y. Z. Zhou12,g, J. Zhu43, K. Zhu1, K. J. Zhu1,58,63, K. S. Zhu12,g, L. Zhu34, L. X. Zhu63, S. H. Zhu70, S. Q. Zhu42, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu71,58, Z. A. Zhu1,63, J. H. Zou1, J. Zu71,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
64 University of Groningen, NL-9747 AA Groningen, The Netherlands
65 University of Hawaii, Honolulu, Hawaii 96822, USA
66 University of Jinan, Jinan 250022, People’s Republic of China
67 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
68 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
69 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
70 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
71 University of Science and Technology of China, Hefei 230026, People’s Republic of China
72 University of South China, Hengyang 421001, People’s Republic of China
73 University of the Punjab, Lahore-54590, Pakistan
74 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
75 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
76 Wuhan University, Wuhan 430072, People’s Republic of China
77 Yantai University, Yantai 264005, People’s Republic of China
78 Yunnan University, Kunming 650500, People’s Republic of China
79 Zhejiang University, Hangzhou 310027, People’s Republic of China
80 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
(February 13, 2025)
Abstract
Based on 368.5 of collision data collected at center-of-mass energies 4.914 and 4.946 GeV
by the BESIII detector, the process
is searched for the first time. No significant signal is observed and
the upper limits at the 90% confidence level on the product of the Born cross section
and the branching fraction at 4.914 and 4.946 GeV
are set to be 0.85 and 0.96 pb, respectively. These measurements provide useful
information for the production of the at colliders
and deepen our understanding about the nature of this particle.
The quark model categorizes hadrons into two types: mesons with a quark and an antiquark, and baryons with three quarks.
Within the framework of the theory of strong interactions, quantum chromodynamics (QCD),
it also allows the existence of more complex structures, generically called exotic hadrons.
A series of exotic hadron candidates, which cannot be accommodated by the potential model Eichten et al. (1978, 1980); Godfrey and Isgur (1985),
were observed experimentally in the charmonium energy region during the past decades.
They are suggested as good candidates
of molecule, quark-gluon hybrid, or tetraquark states Olsen et al. (2018). The well-known state was first observed
by the Belle experiment in the decay Choi et al. (2003),
and confirmed subsequently by several other experiments Acosta et al. (2004); Abazov et al. (2004); Aubert et al. (2005); Aaij et al. (2012); Chatrchyan et al. (2013); Ablikim et al. (2014a).
Ten years after its discovery, its spin-parity quantum numbers
were finally determined to be by the LHCb Collaboration Aaij et al. (2013).
The mass and width are determined to be MeV
and = 1.19 0.21 MeV using a Breit-Wigner resonance model Workman et al. (2022).
Since the discovery of the ,
there have been tremendous efforts to understand its inner structure.
Experimentally, there are intensive studies on the decays currently. The decays of Choi et al. (2003); Ablikim et al. (2014a); Aaij et al. (2020),
Ablikim et al. (2020a); Aaij et al. (2014); Aubert et al. (2006); Bhardwaj et al. (2011), Ablikim et al. (2019a),
Ablikim et al. (2019b); del Amo Sanchez et al. (2010),
Aubert et al. (2008); Aushev et al. (2010); Ablikim et al. (2020a)
have been well observed.
Theoretically, is interpreted as a good candidate of a meson molecule Swanson (2006); Guo et al. (2018)
since its mass is quite near mass threshold.
On the other hand, its quantum number is .
So far, the -wave excited charmonium state (with ) is still missing,
and the might be a good candidate for since its mass is similar to the
potential model prediction Eichten et al. (1978). Other interpretations such as a tetraquark candidate Maiani et al. (2005); Esposito et al. (2015) is also possible.
However, there is still no solid conclusion for the nature of .
In complement with decay, the production of offers a
new window to understand its nature. In 2014, the BESIII experiment observed the
production Ablikim et al. (2014a).
What is intriguing is that the might originate from the radiative transition of an excited vector state , which for the first time brings together two
charmoniumlike states and hints commonality for their underlying nature Olsen et al. (2018). Recently, the
process was observed at BESIII Ablikim et al. (2023a), and the production cross section shows potential enhancement near 4.75 GeV.
In analogy to and , the vector meson
has the same and isospin quantum number. Therefore, the process
of is expected to naturally exist.
By investigating the relative production ratio
Ablikim et al. (2023b)
and also comparing to
Ablikim et al. (2023a, 2024),
we gain a deeper understanding of the production and probe
the potential core component in the wave function Takizawa and Takeuchi (2013). In addition, the decay has also been observed Sirunyan et al. (2020) and has a production rate
Workman et al. (2022).
Together with the process, a more comprehensive understanding
of production will be achieved Maiani et al. (2020).
In this article, we search for the process
using 368.5 pb-1 of data Ablikim et al. (2022) collected with the BESIII
detector Ablikim et al. (2010) operated at the BEPCII storage ring Yu et al. (2016).
The meson is reconstructed via decays, while the is found using decays
with and ( with close branching fraction Workman et al. (2022)).
Due to the mass threshold of the system, only the data at the center-of-mass (c.m.) energies
and 4.946 GeV is used.
The BESIII detector Ablikim et al. (2010) records symmetric collisions
provided by the BEPCII storage ring Yu et al. (2016)
in the center-of-mass energy range from 2.0 to 4.95 GeV,
with a peak luminosity of
achieved at .
BESIII has collected large data samples in this energy
region Ablikim et al. (2020b); Lu et al. (2020). The cylindrical core of the BESIII detector
covers 93 of the full solid angle and consists of a helium-based
multilayer drift chamber (MDC), a plastic scintillator time-of-flight
system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC),
which are all enclosed in a superconducting solenoidal magnet
providing a 1.0 T magnetic field.
A muon chamber (MUC) based on resistive plate chambers
with 2 cm position resolution provides information for
muon identification.
The acceptance of charged particles and photons is 93 over solid angle.
The charged-particle momentum resolution at is , and the
resolution is for electrons
from Bhabha scattering. The EMC measures photon energies with a
resolution of () at GeV in the barrel (end cap)
region. The time resolution in the TOF barrel region is 68 ps, while
that in the end cap region is 60 ps Li et al. (2017); Guo et al. (2017); Cao et al. (2020).
Monte Carlo (MC) samples, simulated using geant4-based software, are used to
optimize the selection criteria, determine the detection efficiency and study the potential backgroundsAgostinelli et al. (2003).
In the BESIII software framework, kkmc Jadach et al. (2001) is the generator used to generate charmonium states by including
initial state radiation (ISR) effects and the spread of the beam energy.
We generate 50000 signal MC events of
at each c.m. energy with a phase space (PHSP) model describing
the uniform angular distribution of the final states.
The
decay is described with the PHSP model.
Final state radiation of charged particles are simulated with the photos
package Golonka and Was (2006). For the possible ISR effect,
we model the -dependent production
cross section with a two-body phase space
is the mass of Workman et al. (2022)). Inclusive MC samples,
with a luminosity which are ten times larger than the data sample,
are generated at each c.m. energy to study the possible backgrounds.
The inclusive MC sample includes the production of open charm
processes, the ISR production of vector charmonium(like) states,
and the continuum processes incorporated in kkmc. All particle decays are modeled with evtgen Lange (2001); Ping (2008) using branching fractions
taken from the
Particle Data Group Workman et al. (2022), when available, and
the remaining unknown charmonium decays are modeled with lundcharm Chen et al. (2000); Yang et al. (2014).
Charged tracks detected in the MDC are required to be within a polar angle ()
range of , where is defined with respect to the -axis,
which is the symmetry axis of the MDC. For charged tracks, the distance of closest
approach to the interaction point (IP)
must be less than 10 cm along the -axis, , and less than 1 cm
in the transverse plane, . For each candidate event, the pions from
decay (kaons from decay) and the leptons from
decay are kinematically well separated. Charged tracks with momenta larger than
1.0 GeV/ in the lab frame are assumed to be leptons, and the others are assumed to be pions or kaons.
The energy deposited in the EMC is used to separate electrons from muons. For both
muon candidates, the deposited energy in the EMC must be less than 0.4 GeV;
while for both electrons, it must be larger than 0.8 GeV. To separate pions
from kaons, particle identification (PID) combines measurements of the specific ionization
energy loss in the MDC (d/d) and the flight time in the TOF to form
likelihoods for each hadron hypothesis.
Charged kaons (pions) are identified by comparing the likelihoods for the kaon (pion) hypotheses with ().
For the candidate events
with detected,
referred to as 6-track events, the
net charge is required to be zero. A four-constraint (4C) kinematic
fit imposing energy-momentum conservation is applied on the 6-track events
to improve resolution and suppress backgrounds. The kinematic fit is required to
be less than 150. The selection criteria are optimized
by maximizing the figure-of-merit
where is the detection efficiency from signal MC events,
is the assumed significance value which is set to 3 and is the expected number of background
events obtained from inclusive MC samples.
To select the resonance, a mass window is
defined as GeV (mass resolution is 6 MeV),
which roughly covers about of the signal.
The signal window of the resonance is set as
GeV (mass resolution is 7 MeV) with barely any efficiency loss according to signal MC events.
To estimate the non- background,
the sideband regions are defined as [3.010,3.065] and GeV,
which are twice as wide as the signal region.
Due to the constraint of the mass threshold of double kaons,
the sideband is defined as GeV, which is
as wide as the signal region.
After applying all of the selection criteria, there is no
background in the 6-track events, as indicated by the inclusive MC sample.
Figure 1 shows the invariant mass distributions of the lepton and kaon pairs
for 6-track events from the full dataset.
Figure 1: The distributions of (a) and (b) for 6-track events.
Dots with error bars are 4.914 and 4.946 data, the red histograms are
the signal MC sample, and the blue filled histograms are the inclusive MC sample.
In order to further improve the detection efficiency and thus the signal yield,
candidate events with
or detected are also reconstructed, with one of the soft kaons
missing due to an inefficiency.
A one-constraint (1C) kinematic fit is performed to these 5-track events,
by constraining the mass of the missing particle to the nominal mass of the
kaon Workman et al. (2022).
The kinematic fit is required to be less than 20,
which is optimized by maximizing the FOM defined in Eq. (Search for the process at BESIII).
Due to the absence of one track in this case, the mass resolution of the event is slightly worse
compared to that of the 6-track events. So
the mass window is defined as GeV (mass resolution is 10 MeV) and
the sideband regions are [3.000,3.065] and GeV
for the 5-track events, which are
twice as wide as the signal region. The signal region
of the resonance is set to be GeV (mass resolution is 8 MeV) and
the sideband is defined as GeV, which is
as wide as the signal region.
There is a case, about of the total signal events, that are detected in
the final state, owing to one of the soft kaon decays in the
detector. This case is classified as 5-track events and reconstructed by missing a kaon.
By analyzing the inclusve MC samples, we find there are some proton backgrounds remaining in the 5-track events.
To reduce the misidentification background in the channel, the MUC is used
to identify muons. At least one of the muon candidates should have a hit depth 30 cm in the MUC.
Figure 2 shows the and distributions for 5-track events from the full data samples,
where is the recoil mass from the system,
and denoted the four-momenta of the initial colliding beams and the system.
Figure 2: The distributions of (a) and (b)
for 5-track events.
Dots with error bars are 4.914 and 4.946 GeV data, the red histograms are the signal MC sample, and
the blue filled histograms are the inclusive MC sample.
After applying all the selection criteria, Fig. 3 shows
the distribution from 6- and 5-track events at each c.m. energy. The invariant mass = is defined, which
partly helps to cancel the resolution effect of the lepton pairs.
Here is the nominal mass of from the PDG Workman et al. (2022). Through studying the inclusive MC samples at each c.m. energy,
no dominant background survives for both 6- and 5-tracks events.
The study of the and mass sideband events also shows the background level is low and so we neglect the
background.
Figure 3: The distributions of from (a) 6-track events at , (b) 5-track events at ,
(c) 6-track events at and (d) 5-track events at , respectively.
Dots with error bars are data, the red histograms are the signal MC sample,
the blue filled histograms are the inclusive MC sample and
the green filled histograms are the 2-dimensional sideband.
The product of the Born cross section of and at c.m. energy is calculated with
where is the number of signal events,
is the integrated luminosity, is the selection efficiency, is
the vacuum polarization factor taken from QED calculation with an accuracy of 0.05 Actis et al. (2010),
is a product of the branching fractions
and , and is the radiative correction factor
calculated by kkmc Jadach et al. (2001).
Due to no events observed in the signal region, we report an upper limit on the Born cross section
at the 90 confidence level (CL) using a frequentist method with an unbounded profile likelihood treatment Rolke et al. (2005).
The number of signal events is determined by the signal region for 6- and 5-track events,
which is defined as GeV.
The possible background estimated by the sideband regions
and GeV from 6- and 5-track events is subtracted
(at the sideband region is defined as GeV
due to the limitation of kinematics).
Assuming the background follows a Poisson distribution and
the efficiency (sum of 6- and 5-track events) follows a Gaussian
distribution with a standard deviation equal to the systematic uncertainty,
the upper limit on the product of the Born cross section of and
the branching fraction of
at the 90 CL at each c.m. energy
are measured and listed in Table 1.
Table 1: The upper limit on the product of the Born cross section
of and the branching fraction
of [denoted as (pb)] at 90 CL
at each c.m. energy.
(GeV) is the c.m. energy, () is the integrated luminosity,
is the number of observed events in signal
region, is the number of observed events in the sideband region,
is the upper limit on the number of observed signal at the 90 CL,
) and ) are detection efficiencies
for the 5- and 6-track events, respectively,
is the radiative correction factor.
4.914
208.11
0
1
1.70
0.690
19.7
2.8
0.85
4.946
160.37
0
0
2.00
0.755
20.8
7.0
0.96
In the cross section measurement, the systematic uncertainties are mainly
from the luminosity measurement, branching fractions, tracking efficiency, PID efficiency,
MUC hit depth requirement, kinematic fit,
radiative correction factor , MC decay model and mass window.
The uncertainty from the luminosity measurement is estimated to be less than 0.66 using
large angle Bhabha scattering events Ablikim et al. (2022).
The uncertainties of the decay branching fractions are quoted from the PDG Workman et al. (2022).
The uncertainty of the tracking efficiency for high momentum leptons is assigned to be 1 per track according
to the study of at BESIII Ablikim et al. (2017).
In this measurement, both one-
and two kaon events are reconstructed.
The total uncertainty of the kaon tracking efficiency
is 1.0(2.0) for 5-track(6-track) events.
Considering the uncertainty of PID efficiency is per kaon track at BESIII,
the kaon PID uncertainty is 1.0(2.0) for 5-track(6-track) events, too.
The systematic uncertainty from pion tracking and PID
efficiencies are both 1.0 per pion track Ablikim et al. (2014b).
The uncertainty of the MUC hit depth is studied using the control sample of Ablikim et al. (2023b). The
difference in efficiency between data and MC simulation due to
the requirement of the hit depth in the MUC is taken as the systematic uncertainty.
A helix parameters correction method is used to estimate the difference between data and signal MC events
caused by the kinematic fit. The difference in efficiency with and
without correction is taken as the systematic uncertainty.
The systematic uncertainty of the radiative correction factor is studied
by comparing the difference between factors obtained with the two-body phase space model and with a flat cross section line shape.
The difference in is taken as the uncertainty.
To estimate the uncertainty due
to the MC model, the angular distribution of
is modeled by a 1 cos
distribution, where the efficiency difference with respect to PHSP is taken as the systematic uncertainty.
The control sample with Ablikim et al. (2021) is selected
to study the system uncertainty caused by the mass window.
The difference in efficiency between data and signal MC events due to the mass window is taken as the
systematic uncertainty.
In the measurement, two subdata samples, i.e. the 6- and 5-track events
are reconstructed. The same source of systematic uncertainties contribute to the two
subdata samples, namely the 5- and 6-track events. These are combined
by taking the weighted average of their systematic uncertainties: Eq. (Search for the process at BESIII) etc.
where is the average systematic uncertainty on the cross section,
while , and are the weight, efficiency and systematic uncertainty
for the th subdata sample, respectively. For the soft kaon decay events (),
its weight contributes to the 5-track subdata sample for the mass window, the MUC hit depth requirement and the kinematic fit;
otherwise its weight contributes to the 6-track subdata sample.
Assuming all these sources are independent, the total systematic uncertainty in the
cross section measurement is obtained by adding them in quadrature.
Table 2 and Table 3 summarize all of the
systematic sources and their contributions at 4.914 GeV and 4.946 GeV, respectively.
Table 2: Systematic uncertainties () in the measurement of the Born cross section at 4.914 GeV.
Uncertainty
5-track
6-track
Weighted average
Luminosity
0.7
0.7
Tracking
5.0
6.0
5.2
PID
3.0
4.0
3.2
0.6
0.6
1.0
1.0
Radiative correction
0.7
0.7
mass window
0.1
0.1
0.1
MUC hit depth
2.1
0
1.8
Kinematic fit
0.7
0.4
0.7
MC model
5.2
5.9
5.4
Total
8.5
Table 3: Systematic uncertainties () in the measurement of the Born cross section at 4.946 GeV.
Uncertainty
5-track
6-track
Weighted average
Luminosity
0.7
0.7
Tracking
5.0
6.0
5.3
PID
3.0
4.0
3.3
0.6
0.6
1.0
1.0
Radiative correction
1.9
1.9
mass window
0.1
0.1
0.1
MUC hit depth
2.9
0
2.2
Kinematic fit
0.6
0.3
0.5
MC model
2.1
14.0
6.2
Total
9.4
In summary, with a data sample corresponding to an integrated
luminosity of 368.5 collected by the BESIII detector, the process of
is searched for the first time. No significant signal is observed and the
upper limits at the 90% CL on the product of the Born cross section
and the branching fraction at 4.914 and 4.946 GeV
are set to be 0.85 and 0.96 pb, respectively. Considering the
cross section pb near the production threshold Ablikim et al. (2023b),
we obtain a rough estimation for the production ratio 9.
It is in the same order as the relative production ratio
Ablikim et al. (2023a, 2024).
These measurements provide important inputs to the production of at colliders, and help constrain the possible component in the
wave function.
With the upgrade of the BEPCII Ablikim et al. (2020b) project, more data in this energy region is expected and a more comprehensive study
of the production will be achieved, which will hopefully reveal the nature of state.
Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts No. 2020YFA0406300, No. 2020YFA0406400, No. 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts No. 11635010, No. 11735014, No. 11835012, No. 11935015, No. 11935016, No. 11935018, No. 11961141012, No. 12025502, No. 12035009, No. 12035013, No. 12061131003, No. 12192260, No. 12192261, No. 12192262, No. 12192263, No. 12192264, No. 12192265, No. 12221005, No. 12225509, No. 12235017;
the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; CAS Key Research Program of Frontier Sciences under Contracts No. QYZDJ-SSW-SLH003, No. QYZDJ-SSW-SLH040; 100 Talents Program of CAS;
Project No. ZR2022JQ02 supported by Shandong Provincial
Natural Science Foundation; supported by the China
Postdoctoral Science Foundation under Grant
No. 2023M742100;
The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; European Union’s Horizon 2020 research and innovation programme under
Marie Sklodowska-Curie grant agreement under Contract No. 894790; German Research Foundation DFG under Contracts Nos. 455635585, Collaborative Research Center CRC 1044, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
Chatrchyan et al. (2013)S. Chatrchyan et al. (CMS Collaboration), “Measurement of the (3872) Production Cross Section Via Decays to in collisions at = 7 TeV,” JHEP 04, 154 (2013), arXiv:1302.3968 [hep-ex] .
Cao et al. (2020)P. Cao et al., “Design and construction of the new BESIII endcap Time-of-Flight system with MRPC Technology,” Nucl. Instrum. Meth. A 953, 163053 (2020).
Chen et al. (2000)J. C. Chen, G. S. Huang, X. R. Qi, D. H. Zhang, and Y. S. Zhu, “Event generator for and decay,” Phys. Rev. D 62, 034003 (2000).
Yang et al. (2014)R. L. Yang, R. G. Ping, and H. Chen, “Tuning and Validation of the Lundcharm Model with Decays,” Chin. Phys. Lett. 31, 061301 (2014).
Actis et al. (2010)S. Actis et al. (Working Group on Radiative Corrections, Monte Carlo Generators for Low Energies), “Quest for precision in hadronic cross sections at low energy: Monte Carlo tools vs. experimental data,” Eur. Phys. J. C 66, 585–686 (2010), arXiv:0912.0749 [hep-ph] .